Martha s preference over two goods x and y is represented by the utility function U = 5xy. Her budget constraint is given by 5x + y = 30

a) Find the MRS

b)Find MRT

c)Given her budget how much of each goods should she consume in order to maximise her utility? show your work.


1
Expert's answer
2022-03-22T11:06:09-0400

A).

U=5xy

budget line;

5x+y=30

MRSxy="\frac{MUx}{MUy}" ="\frac{px}{Py}" ,from the budget line px=$5,Py=$1

="\frac{5}{1}"

MRSxy=5


B).

U=5xy

budget line;

5x+y=30

MRTxy="\frac{MCx}{MCy}" ="\frac{px}{Py}" ,from the budget line px=$5,Py=$1

="\frac{5}{1}"

MRTxy=5


C).

At optimal utlity :

"\frac{MUx}{MUy}" ="\frac{Px}{Py}" ,from budget line Px=$5,Py=$1


MUX="\frac{du}{dx}" 5xy


MUX=5Y


MUY="\frac{du}{dy}" 5XY

MUY=5X

Therefore,

"\frac{5Y}{5X}=\frac{5}{1}"


"\frac{Y}{X}=5"


Y=5X (Substitute in the budget line)


5x+y=30

5x+(5x)=30

10x=30

x=3


5(3)+y=30

15+y=30

y=30-15

y=15

at maximum utility X=3,Y=15